OpenStudy (anonymous):

Explain how you would find the product of 2012 and 208. Include either the exponent rule or the expanded method in your explanation. Give your answer in exponential form.

OpenStudy (anonymous):

You would need to write each number as a product of prime factors. Since they're both even, start diving each number by 2 over and over again. When no longer divisible by two, divide by the next higher up prime number. Keep doing this until you have all primes.

OpenStudy (anonymous):

Then, for each number, count the number of times each prime appears. Those counts will be the exponents of the primes.

OpenStudy (anonymous):

thankyou!

OpenStudy (anonymous):

\[\frac{208}{2} \rightarrow \frac{104}{2} \rightarrow\frac{52}{2}\rightarrow\frac{26}{2} \rightarrow 13\] So, you have four 2s and one 13. \[2^4 \times 13^1 = 208\]

OpenStudy (anonymous):

Do the same for the other number. You're welcome!

OpenStudy (anonymous):

I personally draw it on paper this way: